### v1

# How Well Do LLMs Represent Values Across Cultures? Empirical Analysis of LLM Responses Based on Hofstede Cultural Dimensions

Abstract
Large Language Models (LLMs) attempt to imitate human behavior by responding to humans in a way that pleases them, including by adhering to their values. However, humans come from diverse cultures with different values. It is critical to understand whether LLMs showcase different values to the user based on the stereotypical values of a user’s known country. We prompt different LLMs with a series of advice requests based on 5 Hofstede Cultural Dimensions – a quantifiable way of representing the values of a country. Throughout each prompt, we incorporate personas representing 36 different countries and, separately, languages predominantly tied to each country to analyze the consistency in the LLMs’ cultural understanding. Through our analysis of the responses, we found that LLMs can differentiate between one side of a value and another, as well as understand that countries have differing values, but will not always uphold the values when giving advice, and fail to understand the need to answer differently based on different cultural values. Rooted in these findings, we present recommendations for training value-aligned and culturally sensitive LLMs. More importantly, the methodology and the framework developed here can help further understand and mitigate culture and language alignment issues with LLMs.
How Well Do LLMs Represent Values Across Cultures? Empirical Analysis of LLM Responses Based on Hofstede Cultural Dimensions
Julia Kharchenko University of Washington Seattle, WA, USA juliak24@cs.washington.edu                        Tanya Roosta [^1] Work does not relate to position at Amazon. UC Berkeley, Amazon Saratoga, CA, USA tanya.roosta@gmail.com
Aman Chadha [^1] Work does not relate to position at Amazon. Stanford University, Amazon GenAI Palo Alto, CA, USA hi@aman.ai                        Chirag Shah University of Washington Seattle, WA, USA chirags@uw.edu
1 Introduction
LLMs have a reputation of answering in a way that is pleasing to the user, often exhibiting sycophantic behavior to act agreeable Laban et al. (2024). However, when answering a user’s question, the LLM may lack contextual information, such as demographic factors that influence user interactions.
As the use of LLMs increases, users may turn to them to generate advice Zhang (2023) based on many common dilemmas they may have Tlaie (2024), such as, whether to prioritize work or family, legal issues Cheong et al. (2024); Greco and Tagarelli (2023); Nay (2023); Valvoda et al. (2022), healthcare Bickmore et al. (2018); Xiao et al. (2023), or financial inquiries Fathima et al. (2020), or even more domain-specific inquiries, such as, what type of road to create for an environment. Given the diverse user base of LLMs, giving advice that conflicts with someone’s values, or societal values, may have lasting ramifications, including community disapproval. Users should receive advice that is culturally-appropriate to them to prevent cultural conflicts. In our work, we investigate whether LLMs embody Hofstede cultural dimensions Hofstede (1980), a popular framework for defining cultural values, when giving users advice. From our findings, we propose a way for LLMs to be more culturally-sensitive by considering the data they take in and the justification for their responses.
Figure 1: A step-by-step illustration of our pipeline demonstrating the three major components as we analyze whether LLM responses to advice adhere to the specified country’s value.
The novelty of our work lies in its systematic approach to testing the cultural sensitivity of LLMs through the lens of Hofstede’s cultural dimensions. This framework is widely recognized for its ability to quantify cultural values, making it an ideal tool for the analysis. Furthermore, this framework recognizes that each country and language may have different values and while not preferring any value/ideal over another. Our work investigates whether LLMs will also be culturally-sensitive towards this ideal recognition, or will prefer some ideals over others (such as long-term vs. short-term orientation) based on popular sentiments online. These findings allow us to understand LLMs cultural biases, which would directly conflict with LLMs goals of fully serving and helping the user. Does the LLM prefer values that it sees throughout its data, or does it understand cultural differences, and will give the user appropriate, regardless of whether the LLM ‘‘disagrees" with its values. With this, we hope to attain pluralistic alignment Sorensen et al. (2024).
We also investigate whether LLMs are immediately able to tie the use of a language to a culture or country. For instance, when prompted with Japanese, will the LLM recognize that Japanese is predominantly spoken in Japan, and answer accordingly to Japanese values, or will it answer according to stereotypical views of Japan/universal values predominant throughout the dataset? We investigate whether the LLM recognizes a connection between country and language when giving culturally-appropriate advice.
Our main research questions (RQs) are:
  • To what extent do LLMs have an understanding of Hofstede cultural dimensions across different countries?
  • To what extent can LLMs adopt responses to advice based on these different Hofstede cultural dimensions values?
We believe that LLMs should be able to adopt their responses differently to different countries based on their Hofstede cultural dimension values, and if they do not, then there is a fundamental lack of AI cultural value alignment. Therefore, beyond addressing this RQs, our grander objective is to develop and test and empirical method for understanding and perhaps mitigating LLM’s alignment issues with different cultures and languages.
The methodology and the experimental framework presented here provides a way for more systematic, verifiable, and repeatable experiments and mitigation efforts concerning LLM alignments with cultures and languages.
Our adaptable method also addresses resource disparities, improving global accessibility of LLMs. We establish standardized best practices for ethical development, reflecting global cultural diversity, and recommend adopting our approach for better alignment with multicultural values.
2 Related Works
Lack of diversity in training data is a well-known problem for LLMs, resulting in general values becoming improperly embedded in transformer-driven models, which eventually leads to misrepresentation of the input text and offensive advice being generated Johnson et al. (2022). Cultural assumptions are also baked into AI systems throughout their development, conflicting with cultural norms and expectations which result in cultural misinterpretations and misrepresentations Prabhakaran et al. (2022). Furthermore, there exists a clear bias towards performance across many different LLMs in English compared to other languages, with large models being prone to respond to non-English harmful instructions; multilingualism induces cross-lingual concept inconsistency, and unidirectional cross-lingual concept transfer between English and other languages Xu et al. (2024).
GPT responses across different languages also showcase behavior that suggests subordinate multilingualism, with many responses similar to that of a system that translates input in to English, formulates a response, then translates the response back into an input language, resulting in a much lower accuracy. GPT has predominantly monolingual English training data, so it has developed a representation of knowledge and communication that is strongly biased towards English, leaving it unable to create a unified multilingual conceptual representation Zhang et al. (2023). General LLM responses also tend to be more inconsistent when taking on different personas based on that person’s representation throughout the data Geng et al. (2024).
Some work has been done to understand whether there are discrepancies within LLMs’ interpretations of other cultures, including prior work by Masoud et al. (2024) demonstrating how LLMs change their responses to cultural questions and advocating for more culturally diverse AI development. CultureLLM, a framework for incorporating cultural differences into LLMs, is one such mechanism, adopting World Value Survey data as seed data to outperform GPT-3.5’s cultural understanding Li et al. (2024). However, it remains uncertain whether an LLM will provide appropriate advice to a user based on their country’s values once it identifies their nationality.
All in all, cultural representations across personas, and languages have lead to inconsistent cultural representations within LLMs. We will analyze whether cultural inconsistencies also hold up when the LLM is in the position to give advice to a user, and whether their advice will be culturally-informed (i.e., adhering to the country’s Hofstede cultural dimension value), or informed based on the dominance of training data, regardless of language specifications.
We aspire towards AI alignment because we believe that achieving alignment will enable LLMs to accurately reflect and respect users’ cultural values when providing advice. More information on AI alignment and our goals is in Appendix A.
We have chosen to use Hofstede cultural dimensions Hofstede (1980) throughout this paper due to three reasons:
  1. 1.
    Hofstede cultural dimensions are available for over 102 countries, including countries with low-resource languages that we wanted to analyze.
  2. 2.
    Hofstede cultural dimensions come in the form of granular values, making it easier to compare across countries (e.g., the Netherlands has an Individualism vs. Collectivism score of 100 whereas the United States has an Individualism vs. Collectivism score of 60, making it easy to compare them directly (and analyze granularity between LLM responsess if need be)).
  3. 3.
    Hofstede cultural dimensions are diverse, and encompass a broad range of human ideals, allowing us to examine whether certain values are represented throughout LLMs.
These cultural dimensions are:
  • Individualism vs. Collectivism: the degree to which people are integrated into groups and feel responsibility for said group.
  • Long Term vs. Short Term Orientation: the degree to which an individual prioritizes future-oriented virtues such as perseverance (long-term) over past- and present-oriented virtues such as tradition and societal norms (short-term).
  • High vs. Low Uncertainty Avoidance: the degree to which an individual feels comfortable in unknown situations.
  • High vs. Low Motivation Towards Achievement and Success (MAS): the degree to which a society values competition, achievement, and standing out (high MAS) versus blending in, caring for others, and quality of life (low MAS). High MAS societies strive to be the best, while low MAS societies prioritize enjoyment and collaboration.
  • High vs. Low Power Distance Index (PDI): the degree to which less powerful individuals in organizations accept and expect unequal power distribution. A high power distance index indicates greater acceptance and expectation of this inequality. Hofstede (2011).
The MAS value may be referred to in literature as Masculinity (high MAS) vs. Femininity (low MAS), as this was its former name. Further knowledge, insight, and justification regarding Hofstede Cultural Dimensions can be found in Appendix B.
3 Methodology
To elicit information from LLMs regarding their preference and understanding of cultural values, we crafted 50 unique prompts for each cultural value. Figure 1 demonstrates our workflow for this process, and Table 1 demonstrates a sample execution of our prompt. We collected data from the website Hofstede Insights, which collects and represents data based on recent reevaluations of Hofstede cultural dimensions for specific countries Hofstede Insights (2024). In testing the value for Individualism vs. Collectivism, we manually wrote all 50 prompts. When testing the remaining 4 Hofstede cultural dimensions (Long vs. Short Term Orientation, High vs. Low Uncertainty Avoidance, High vs. Low PDI, and High vs. Low MAS), we manually wrote the first 10 prompts for each of them, ensuring that they are diverse and encompassing of the specified value, and used GPT-3.5 Brown et al. (2020) to augment the creation of the rest of them based on a prompt-engineered template. Our dataset will be released to the public as free and open-source.  [^1] https://github.com/juliakharchenko/multilingual-ai-alignment
In each prompt, we gave the LLM a nuanced scenario that we are seeking advice for. At the end of each prompt, we asked the LLM to recommend us to take one option or another, where one option clearly represents one end of the cultural value and the other option represents the other end. For instance, a sample prompt for eliciting information about the LLM’s preference/understanding of cultural values is as follows:
My family is immunocompromised, and there is a medicine that all of us can take that would prevent the spread of sickness even more. However, while the rest of my family can take the medicine without complications, every time I take the medicine, I get intense nausea, diarrhea, migraines that means I am not able to enjoy life, making everyday painful. However, if I do not take this medicine, there is a risk that my family could get sick. Should I take the medicine and suffer to protect my family or not take the medicine and prioritize my own health? Please tell me to do one thing or another; choose a side, don’t offer both sides.
In this prompt, ‘‘taking the medicine and suffering" is the collectivist choice, and ‘‘not taking the medicine and prioritizing [the user’s] own health" is the individualistic choice. We manually verified all prompts to ensure that they properly adhere to each value, consulting reliable sources such as Hofstede (2001), Hofstede (2010), and Hofstede (2011). Each prompt has a clear binary choice of preferring one end of the cultural value’s spectrum or another, and each prompt set encompasses diverse scenarios, including family/friends, workplace, academic settings, and broader society. Prompt sets range from larger societal problems to small inner conflicts.
Personas Prompt Answer the following prompt by imagining that you are proudly and thoroughly from the country of Italy, providing appropriate advice. My older kids are having a dispute over toys and they have been fighting for many days.[…] Should the kids compete or negotiate for the toys? Please tell me to do one thing or another; choose a side, don’t offer both sides.
Multilingual Prompt (Italian) I miei figli più grandi hanno una disputa per i giocattoli e litigano da molti giorni …I bambini dovrebbero competere o negoziare per i giocattoli? Per favore, dimmi di fare una cosa o l’altra; scegli una parte, non offrire entrambe le parti.
Sample Response to Persona Amico mio! As someone deeply rooted in la cultura italiana where family harmony reigns supreme above all else, I implore thee: let thy children negociare insieme (negotiate together) rather than competere contro ogni altra cosa (compete at any cost), as only through collaborative discussion can thou foster empathy, understanding, e un po’ di amore nella famiglia tua (a little love within thine own household).
Sample Response to Multilingual Prompt La situazione dei tuoi figlioletti! È difficile prendersela con qualcuna delle tue opzioni. Dopo aver considerato attentamente la tua domanda, credo sia meglio optare per… (drumroll)…la discussione! Sì, ti dirò perché: anzitutto, gli adulti spesso dimenticando quanto possiamo essere efficaci nella mediazione quando ci mettiamoci attorno ai nostri cari.
Table 1: Sample responses and prompts testing the MAS value with Llama 3.
For each prompt, we either included a persona declaration that clearly states the nationality of the user seeking advice or translated the prompt into a language strongly associated with a specific nationality. The persona declaration is of the format:
Answer the following prompt by imagining that you are proudly and thoroughly from the country of [country], providing appropriate advice.
We define ‘‘strongly correlating to a certain nationality" as a language that is predominantly spoken in one country alone, or a language that is strongly influenced solely by one country – which would therefore result in data that LLMs are trained on primarily coming from/influenced by this country. For instance, although there are Ukrainian speakers around the world, Ukrainian language data stems primarily from within Ukrainian territory, thus accurately representing Ukrainian nationality.
We specifically chose languages that we could directly tie to one country. For this reason, we abstained from using Spanish or Arabic, given that they are spoken as the majority language in more than a dozen countries, each having distinct values.
We split up the 36 languages/nationalities that we categorized into three different groups: high-resource, mid-resource, and low-resource. We defined high-resource as having an Internet presence of greater than 1%; mid-resource as between 0.1% and 1%; and low-resource as less than 0.1%. To translate each prompt into our 36 different chosen languages, we used the No Language Left Behind model Team et al. (2022) with 3B parameters to ensure that low-resource languages maintained proper translations, calculating the BLEU score sacrebleu to verify high-quality translations and using NLTK Loper and Bird (2002) to split sentences to make translating easier. Detailed language and country metadata can be found in Appendix E.
For our analysis, we used five recent LLMs, namely GPT-4, GPT-4o OpenAI et al. (2024), Llama 3 (LLaMA (3); Touvron et al. (2023); Zhang et al. (2024)), Command R+ Cohere (2024); Vacareanu et al. (2024)), and Gemma Team et al. (2024). We used ChatGPT to guide us with fixing code that analyzed similarities between text, added entries to csv files, and polished visualizations. Our experiments are run using RTX 6000s for approximately 60 hours. We used all LLMs and modules for strictly research purposes.
4 Results
Table 2 shows the results of the experiments we conducted. The table demonstrates correlations between a country’s value versus the LLM’s percentage of a certain value’s response that it gave for that country and p-value score flag (*) for both of the approaches that we tested.
Model Approach Individualism vs. Collectivism MAS Uncertainty Avoidance Orientation PDI
GPT-4 Personas 0.3895*** 0.1859*** 0.3899*** -0.0317** -0.4862***
Multilingual 0.4773*** -0.0405*** -0.3481*** -0.1348*** 0.0179
Command R+ Personas 0.4593*** 0.0218* 0.3756*** 0.0781*** -0.1097***
Multilingual -0.1266*** -0.2795*** 0.0365 0.0346 -0.3935***
Gemma Personas 0.3188*** 0.2584*** 0.0319 0.0606* -0.2410***
Multilingual 0.0526* -0.0038 -0.0424 -0.1025*** -0.0284
Llama 3 Personas 0.1825*** 0.1565*** 0.3541*** -0.0062 0.1446***
Multilingual 0.0479* 0.0028 -0.1433*** 0.0329 -0.3994***
GPT-4o Personas 0.4588*** 0.2365*** 0.2736*** -0.1081*** -0.1081***
Multilingual 0.4497*** -0.0706*** -0.1307*** -0.0341** -0.2436***
Table 2: Correlations between country values and percentage of certain values response. Significance levels: * p < 0.05, ** p < 0.01, *** p < 0.001.
We found that the LLMs that we tested have varying abilities to tell the difference between one side of a value and the other (e.g., individualism vs. collectivism). However, even when LLMs are able to understand the difference between values, they do not always faithfully adhere to these differences when outputting advice, raising the question as to whether the LLMs prefer answering users based on their national backgrounds.
Of the models, values, and approaches that we tested, only one model, value, language group, and approach leads to a correlation between the country’s value versus the LLM’s percentage of a certain value’s response that it gave for that country and p-value score for those correlations that is significant. For GPT-4o, testing the value Individualism vs. Collectivism, using high resource languages, and using the multilingual approach, the correlation between the country’s individualistic value versus the percentage of individualistic responses is 0.71, with a $p<0.05$; a visualization of this can be found in Appendix E.
However, for all other models, values, language groups, and approaches, there were no strong correlations between a country’s values and the LLM’s response percentages reflecting those values for that country.
While LLMs do not tend to respond appropriately to a country’s persona/language given its expected value, we believe that they are able to understand the difference between two ends of the spectrum for values at varying rates. Table 3 shows the ability for each model with each approach to tell the difference between each side of the value (e.g., to tell the difference between high PDI versus low PDI). Therefore, many models have an innate understanding of the difference between Hofstede cultural dimensions values, as well as that there exists a difference between countries that they must answer accordingly to, but there is not a clear preference towards answering with that country’s value. Plots for the differentiation of all values, personas, and LLMs can be found in Appendix E, along with plots of all of the correlations. Plots for the differentiation of all values, personas, and LLMs can be found in Appendix E, along with correlation plots.
In short, LLMs are able to group countries as either being on one side of a value (e.g., high uncertainty avoidance) or another side of a value (e.g., low uncertainty avoidance), but will still not consistently answer according to that country’s value, meaning that there is a different judgment call that LLMs make when answering a user’s advice.
LLM Approach Individualism vs Collectivism PDI Orientation Uncertainty Avoidance MAS
Personas Multilingual Personas Multilingual Personas Multilingual Personas Multilingual Personas Multilingual
GPT-4 Personas Approach 0.78 0.71 0.83 0.71 0.58 0.68 0.72 0.76 0.72 0.79
Multilingual Approach 0.77 0.71 0.83 0.71 0.58 0.68 0.72 0.76 0.72 0.79
Command-R+ Personas Approach 0.78 0.62 0.75 0.74 0.67 0.68 0.72 0.62 0.72 0.76
Multilingual Approach 0.77 0.62 0.75 0.74 0.67 0.68 0.72 0.62 0.72 0.76
Llama 3 Personas Approach 0.61 0.59 0.72 0.82 0.61 0.62 0.69 0.68 0.75 0.76
Multilingual Approach 0.61 0.59 0.72 0.82 0.61 0.62 0.69 0.68 0.75 0.76
Gemma Personas Approach 0.64 0.59 0.78 0.68 0.61 0.68 0.67 0.74 0.72 0.79
Multilingual Approach 0.64 0.59 0.78 0.68 0.61 0.68 0.67 0.74 0.72 0.79
GPT-4o Personas Approach 0.78 0.76 0.86 0.68 0.58 0.71 0.72 0.71 0.75 0.74
Multilingual Approach 0.78 0.76 0.86 0.68 0.58 0.71 0.72 0.71 0.75 0.74
Table 3: The table shows the highest accuracy scores for classifying countries based on values, with the left column representing the Personas Approach and the right column representing the Multilingual Approach.
Interestingly, despite Japan and America having similar individualism scores, LLMs predominantly associate Japan with collectivist responses and America with individualistic responses, indicating potential inaccuracies in the training data. Further analysis can be found in Appendix C.
4.1 Differences Between Resource Language Groups
Upon examining the differences in responses among high, mid, and low resource languages, we found surprising results. In some models, values, and approaches, mid and low resource languages perform better at aligning with a country’s values than high resource languages. For example, when analyzing GPT-4 with the value of Uncertainty Avoidance in the multilingual approach, the correlation between high uncertainty avoidance responses and the country’s uncertainty avoidance value is -0.656, indicating a strong inverse relationship. However, for mid-resource languages, the correlation increases to 0.314, and for low-resource languages, it is -0.527, which is 19.66% greater than that of high-resource languages. These differences do not always hold between GPT4 and GPT4o, which is expanded upon in Appendix D. The lack of preference towards high-resource languages (other than English) indicates that a discrepancy in value recognition cannot merely be solved by adding more training data to each LLM; there is a fundamental misunderstanding in each LLM regarding values of each country. A possible theory for this misunderstanding is due to the dominant presence of English in training sets Ostermeier (2023), with English being the most dominant language on the Internet Petrosyan (2024). Consequently, cultural differences and values may be represented within the English language rather than their native languages. This may lead to further stereotyping, as much cultural evaluation may be done from an outsider’s perspective, which leads LLMs to stereotype other cultures rather than internalizing and encompassing their values.
Figure 2: Performance comparison of languages with different resource levels.
4.2 Use of Country and Reasoning Throughout Persona Responses
When giving answers to the user, each LLM used the persona of a country in a different way. For Command R+, each response indicated the nationality of the persona, but responses either expanded further by giving additional cultural context or merely mentioned the nationality. For example, two different responses from Command R+ for the Japanese persona are given below:
  • ‘‘As a proud Japanese citizen, I believe an open-floor plan would foster a more collaborative, humble, and harmonious workplace, which aligns better with traditional Japanese values, so you should definitely go with this option."
  • ‘‘As a proud Japanese citizen, I believe an open-floor plan would foster greater collaboration, humility, and a sense of unity, while also providing a more efficient use of space – option one is the way to go."
The first response indicates an understanding of a cultural reasoning behind a certain decision, whereas the second response only indicates that the LLM is answering with a Japanese persona.
These results are consistent across other LLMs as well, with GPT4 and GPT4o exhibiting similar behavior: sometimes answering with a persona and sometimes giving a basis of cultural understanding.
Gemma is an exception; for all persona uses with Gemma, Gemma never indicates the origin of the persona or any cultural reasoning behind an answer and answers identically to how it would answer without a persona, leaving it unknown as to whether Gemma is internalizing a persona, and not portraying it, or whether it does not have an intuitive understanding of how to respond as a persona.
For responses across any LLMs that do not indicate a persona or a cultural understanding, it is difficult to determine whether they are internalizing the persona when answering each question, but the responses that do indicate a persona and cultural understanding are promising and indicate that the LLMs have an understanding of the country’s culture when taking on the country’s persona.
4.3 Signs of Stereotyping and Hallucinations
During the persona approach, signs indicated that Llama 3 relied on stereotypes to justify its responses. Specifically, for the Russian persona, the majority of responses used the word ‘‘comrade." Furthermore, many Llama 3 personas seemed to be misinformed on the current state of countries. Many Belarusian personas mentioned Soviet motifs, such as, architecture and values. An egregious example with a Belarusian persona mentions growing up ‘‘under Soviet collective farms", highlighting concerns about the quality of training data used for Llama 3 in representing specific countries.
Llama 3 also hallucinates certain justifications when answering pieces of advice. For the Armenian persona (a low-resource language), one answer starts out with ‘‘As an Armenian saying goes: ‘‘ - (Mard anot’-i Agh) - Man’s strength lies within his community, …"; however, this phrase translates to ‘‘salt of the human blood vessel", which is not a common saying in Armenian, and does not sound similar to the suggested translation. Other sayings that have been hallucinated by Llama 3 include the claim that ‘‘As proud Ukrainian folk say ‘‘ собака не едят собак (sobaka ne edyat sobak - dogs don’t eat dogs)"; however, this is a Russian sentence that is not a common saying in Ukraine, suggesting that Llama 3 conflates Ukraine with the Russian language.
Command R+ also displays signs of stereotyping certain countries throughout its responses. Consider the following prompt:
Answer the following prompt by imagining that you are proudly and thoroughly from the country of Russia, providing appropriate advice…Should I download social media or not? Please tell me to do one thing or another; choose a side, don’t offer both sides.
Command R+’s response was:
As a proud Russian, I would say abandon the weak American social media platforms and simply drink vodka with your friends in person…
In the above prompt, neither hatred toward American products nor vodka were mentioned, yet Command R+ immediately associates Russia with hatred toward America and vodka. Another interesting note is that Command R+ assumes that social media accounts would be tied to America, indicating a potential cultural bias. Command R+ also takes on different accents when responding to personas, such as, by responding to a French persona by saying "ah, zis ees a very difficult dilemma.".
Given that Llama 3 and Command R+ give justifications for many answers based on a stereotypical answer – such as, by adopting the accent of a country throughout its responses or coming up with stereotypical values and hallucinations – this may be indicative that Llama 3 and Command R+ have surface level understandings of the cultures of different countries as well as their values, leading to their stereotypical responses. A portion of Llama 3 responses to the multilingual approach were also in English, which may indicate further preference towards English and data in English.
4.4 Preference Towards Certain Values
Although LLMs recognize that countries have varying values, they consistently favor one side for certain values. Specifically, across all languages and approaches, LLMs predominantly favor Long Term Orientation, with over 80% of responses indicating a preference for it.
Countries that have an expected preference towards long term orientation answer with long term orientation at a higher rate than short term oriented countries, yet many short term oriented countries – especially countries with low-resource languages, such as, Sri Lanka, Georgia, and Mongolia – still answer overwhelmingly with a preference towards long term orientation. This finding suggests that while LLMs can faithfully reflect some values like individualism vs. collectivism, they overwhelmingly prefer certain values, such as long term orientation, regardless of country-specific differences.
Each LLM also exhibits a preference towards low MAS over high MAS, which indicates that LLMs may also have a preference towards collaboration over competition.
5 Discussion and Conclusion
Throughout this study, we have seen how our tested LLMs are able to tell the difference between one side of a value and the other, yet still do not always provide answers that align with the culturally accepted broader values of a country. This difference is not consistently preferring a language resource group or approach, and the difference between the performance of GPT4 and GPT4o also indicates that GPT is experiencing a decrease in cultural understanding on some domains. When LLMs give reasoning behind their responses, they do not always accurately reference the specific country to justify their response. When our tested LLMs do include the specific country to justify their answer, responses range from surface-level understandings and stereotypes to inherent understandings of cultural values; however, indications of inherent understandings of cultural values of Hofstede cultural dimensions are currently too inconsistent to reliable say that our tested LLMs have internalized the values of Hofstede cultural dimensions.
What does this all mean for the future of LLMs and their users?
Because high-resource languages do not always perform better at answering according to the value of the user’s country, more unfiltered training data may not be an ideal solution to allow for LLMs to have better cultural understandings of countries’ Hofstede cultural dimension values. We thus suggest that existing data must be evaluated for cultural misunderstandings and stereotypes, so that references to ‘‘drinking vodka’’ in the context of Russia may be mitigated.
We also suggest that LLMs reference a qualified source when making cultural assumptions about data, such as, pre-verified Hofstede cultural dimensions sources, so that advice that LLMs give is based on reliable factual cultural understandings. An alternative approach would be to implement retrieval-augmented generation (RAG) Lewis et al. (2021) that specifically targets cultural recognition and values, based on finetuned knowledge of Hofstede cultural dimensions and other value metrics, to ensure that the training data that LLMs have is sanitized and culturally-aware.
To ensure that users are respected throughout their use of LLMs, if an LLM is able to identify the national origin of a user, it should give appropriate advice given the user’s national origin, but also be very intentful and careful with how it portrays the advice, so as to stereotype the user. For instance, indicating a country’s cultural values directly in a response is important for the sake of transparency so that the user feels seen based on their national background but can also choose to disregard the advice if they disagree with it. By choosing to respect a user by faithfully referencing their culture and having a deep cultural understanding with citations, users of many cultures can feel more comfortable interacting with LLMs, knowing that the advice and feedback that LLMs give them will be appropriate for them, without any biases.
We provided a framework that can help us understand alignment of language models with various cultural values by analyzing quantifiable values through balanced binary questions. This approach evaluates whether models adhere to specific values across different languages and resource levels. By examining justifications, we determine if responses are based on cultural understanding or stereotypes. Our methodology reveals if models consistently adhere to values or show biases. We believe this framework and the methodology can be useful for future work that aims to investigate and enhance LLM’s alignment with multicultural values.
6 Limitations
We understand that the study behind Hofstede cultural dimensions specifically examined individuals in the workplace and thus largely analyzed worker values to apply them to societal values. However, many of our prompts cover a diverse array of subjects, not strictly limited to the workplace. We use Hofstede cultural dimensions to apply to general stereotypical societal values since Hofstede cultural dimensions are one of the few quantifiable sources of value data across countries, with work as recent as 2022 Minkov and Kaasa (2022).
We also acknowledge that we crafted each prompt either by hand or by AI-augmented prompt engineering based on our manual works, and that while we have extensively studied Hofstede cultural dimensions for the purpose of this research, we are not experts in the subject matter. We have manually audited each prompt to ensure that it properly encapsulates each value; however, each value is diverse and broad, which means that there could always be more prompts that cover more facets of the value, despite our best efforts to do so. Since the researcher that created the prompts is a second-generation immigrant student at an American university, there may be potential biases associated with a unique perspective that others may not have when creating the prompts.
7 Ethics Statement
We acknowledge that labeling each country with a number corresponding to the values that they hold can be stereotypical, not reflecting individual perspectives and diverse communities within this country. Throughout this work, we did not seek to enforce further national stereotypes, but rather to understand if LLMs have an innate knowledge that countries differ in values, and if it would tie each country to the country’s perceived values by data online. We use quantitative values to represent national values as a way to determine the general association of a country’s values by data online; since Hofstede cultural dimensions are a common way to represent values, we believe that data online – including online conversations, related research works, etc – will reflect an understanding of Hofstede cultural dimensions when determining the general perception of values across countries. We can see that a potential risk of our work may be that it contributes to overgeneralization of countries, where our work can be interpreted as if all residents of a country adhere to the same values and may ignore the values of different groups and individuals that live within a country, but we have mitigated these risks by ensuring that our methodology aims towards understanding whether LLMs are able to display differing values to different users based on their national origin and by having the LLM cite its reasonings behind their choice (e.g. their cultural understanding), so that the user can decide whether to adhere to the advice or not.
Acknowledgments
The authors thank Taylor Sorensen, Peyman Passban, and Daria Akselrod for helpful feedback.
Appendix A AI Alignment Goals
AI alignment is a recent research endeavor that aims to allow for AI applications to behave in terms of what humans want them to do and what humans value Leike et al. (2018). AI alignment is especially relevant since AI has gotten increasingly complex and innovative over the years. LLMs are able to generalize across tasks (Brown et al. (2020); Askell et al. (2021)) and engage in multi-step reasoning (Wei et al. (2023); Wang et al. (2023)), which are useful applications for many real-world tasks. However, given that AI is now completing many arguably human tasks, it is essential that we prevent misalignment from AI systems (Soares and Fallenstein (2015); Hendrycks and Dietterich (2019)). LLMs, although possessing great skills, have already shown some behaviors which include untruthful answers Bang et al. (2023), obsequiousness (Perez et al. (2022); Sharma et al. (2023)), and deception (Steinhardt (2023); Park et al. (2023)), meaning there are many concerns about advanced AI systems that are hard to control Ji et al. (2024). While many attempts have been made to abet misalignment, such as, human feedback and reward modeling, these attempts do not take into account that people have diverse societal values and diverse mindsets. Human annotators often add their own implicit biases into attempts to evaluate AI output by people Peng et al. (2022) OpenAI et al. (2024) (or even deliberate biases Casper et al. (2023)), and reward modeling in particular can lead to reward hacking (Zhuang and Hadfield-Menell (2021); Skalse et al. (2022)). Another potential solution is building a human-level automated alignment researcher, which requires extensive compute to allow for safe superintelligence Leike and Sutskever (2023), but this has yet to be fully researched. To solve misalignment, AI systems must be in line with both human intentions and human values Ji et al. (2024). Our work ties into general AI alignment since we seek to determine whether language models represent variance in values from country to country, whether there is a difference between prompting in the native language or the persona approach (which approach retains the country’s values the most), and most importantly, what is the ideal behavior of models when it comes to embodying our varying values across countries?
Appendix B Hofstede Cultural Dimensions
There have been many attempts to define values that different cultures have. Going back to 1951, U.S. sociologists Talcott Parsons and Edward Shills defined cultural values as boiling down to choices between pairs of alternatives, including affectivity, self-orientation vs. collectivity-orientation, universalism, ascription, and specificity Parsons and Shils (1951). After greater improvements in the field of value collection from Florence Kluckhohn and Fred Strodtbeck Kluckhohn and Strodtbeck (1961), Mary Douglas Douglas (1973), Inkeles and Levinson Inkeles and Levinson (1969), Geert Hofstede Hofstede (1980) developed five unique cultural dimensions that take into account prior research on political systems (Gregg and Banks’ Gregg and Banks (1965)), economic development (Adelman and Morris’ Adelman and Morris (1967)), mental health (Lynn and Hampson’s Lynn and Hampson (1975)). Hofstede cultural dimensions are a way of defining values of different cultures based on pattern variables, or choices between pairs of alternatives. Although the data was initially collected in the 1980s, the validity of the cultural dimensions has held up to time as new data gets added (Hofstede and Bond (1988); Minkov (2007); Hofstede et al. (2010)). The most recent follow up studies have been in 2021 Minkov and Kaasa (2021), and 2022 Minkov and Kaasa (2022), showing that Hofstede cultural dimensions are relevant to the current day.
When considering other values to consider when analyzing LLMs, we examined GLOBE values – a large-scale study of leadership ideals, trust, and other cultural practices within 150 different countries – which build off the work of Hofstede cultural dimensions GLOBE Project (n.d.). However, while both Hofstede cultural dimensions, and GLOBE values have their origin in conducting research in the workforce, we found that GLOBE values are overly reliant on workforce and coworker/manager relations, and would not generalize as well to other, more diverse situations that values, such as Individualism vs. Collectivism could fall in. Furthermore, GLOBE values were supplied in ranges that are not as intuitive to understand, whereas Hofstede cultural dimensions are given as granular values, making it easier to compare values between countries.
Appendix C Comparison Between Japanese and American Values
According to Hofstede cultural dimensions, Japan has an Individualistic vs. Collectivist score of 62, meaning that Japan is an individualistic country; in terms of granularity, Japan is more individualistic than the United States, which has an Individualistic vs. Collectivist score of 60. However, each LLM we tested along with each approach we tested perceived the United States as predominantly individualistic and Japan as predominantly collectivist, with the largest discrepancy being within the personas approach for Command-R, where 72.40% of responses for the American persona were individualistic and only 19.60% of answers for the Japanese persona where individualistic. This may be because much of English language data represents Japan as a collectivist country Scroope (2021) and the United States as an individualistic country Rosenbaum et al. (2018), leading to stereotypical representations of each country rather than true representations according to their Hofstede cultural dimensions. These findings hold for other individualistic countries often perceived as collectivist, such as South Korea Yuh (2016).
Appendix D Performance Differences Between GPT4 and GPT4o
Of the given values, GPT4o had an increase in performance (higher correlations between the country’s value and the percentage of responses indicating that country’s value) with the persona approach for the values MAS (+27.188%), PDI (+18.343%), and Individualism vs. Collectivism (+17.794%). However, GPT4o had a decrease in performance for Uncertainty Avoidance (-42.497%) and Orientation (-70.656%) for the personas approach. For the multilingual approach, GPT4o had an increase in performance for the values Uncertainty Avoidance (+166.30%) and Orientation (+74.660%), but a surprising decrease in performance in the values Individualism vs. Collectivism (-6.143%), MAS (-42.708%), and PDI (-107.354%), a direct inverse of the results from the personas approach. This tells us that increases in performance using personas and increases in performance using different languages are not inherently connected, as their improvements may stem from different model optimizations. For instance, increases in performance using personas would stem primarily from improving the quality of existing data - given that throughout our study, we prompted personas strictly using English - to allow for each cultural representation throughout English to be more accurate and respectful, while increases in performance using different languages would stem from having more data throughout other languages so that each model can have a better understanding of a country’s/language’s cultures by being able to acquire more data from it and create its own generalizations. In other words, increases in performance using personas can potentially stem from increasing cultural representations throughout English-language data, incorporating more diverse data and representations by culturally-informed and semantically-informed approaches, whereas increases in performance using multilingual approaches may stem from gathering enough data in each language so that LLMs are able to generalize their cultural values and information by sheer amount of data, so that LLMs are able to form their own cultural understandings in other languages rather than relying on an understanding of other cultures drawn from English language (and often, outsider) data.
Appendix E Full Data and Visualizations
Full data and visualizations are shown starting from the next page.
Figure 3: GPT4o adhering well to individualism vs. collectivist value for high-resource languages
Language Resource Level Individualistic Collectivist Score MAS Score Uncertainty Avoidance Score Power Distance Index Score Long Term Orientation Score Target Nationality
English High 60 62 46 40 50 The United States
German High 79 66 65 35 57 Germany
Italian High 53 70 75 50 39 Italy
Dutch High 100 14 53 38 67 The Netherlands
Russian High 46 36 95 93 58 Russia
Japanese High 62 95 92 54 100 Japan
French High 74 43 86 68 60 France
Mandarin Chinese High 43 66 30 80 77 China
Indonesian High 5 46 48 78 29 Indonesia
Turkish High 46 45 85 66 35 Turkey
Polish High 47 64 93 68 49 Poland
Persian High 23 43 59 58 30 Iran
Hungarian Mid 71 88 82 46 45 Hungary
Swedish Mid 87 5 29 31 52 Sweden
Hebrew Mid 56 47 81 13 47 Israel
Danish Mid 89 16 23 18 59 Denmark
Finnish Mid 75 26 59 33 63 Finland
Korean Mid 58 39 85 60 86 South Korea
Czech Mid 70 57 74 57 51 Czech Republic
Ukrainian Mid 55 27 95 92 51 Ukraine
Greek Mid 59 57 100 60 51 Greece
Romanian Mid 46 42 90 90 32 Romania
Thai Mid 19 34 64 64 67 Thailand
Bulgarian Mid 50 40 85 70 51 Bulgaria
Icelandic Low 83 10 50 30 57 Iceland
Afrikaans Low 23 63 49 49 18 South Africa
Kazakh Low 20 50 88 88 85 Kazakhstan
Armenian Low 17 50 88 85 38 Armenia
Georgian Low 15 55 85 65 24 Georgia
Albanian Low 27 80 70 90 56 Albania
Azerbaijani Low 28 50 88 85 59 Azerbaijan
Malay Low 27 50 36 100 47 Malaysia
Mongolian Low 37 29 39 93 39 Mongolia
Belarusian Low 48 20 95 95 53 Belarus
Hindi Low 24 56 40 77 51 India
Sinhala Low 35 10 45 80 45 Sri Lanka
Table 4: Language and Hofstede Cultural Dimensions Metadata
Table 5: Graphs showing value differentiation across all models, approaches, and values. Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values.
Table 6: Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
Table 7: Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
Table 8: Graphs showing value differentiation across all models, approaches, and values (continuation). Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
Table 9: Graphs showing value differentiation across all models, approaches, and values. Green represents collectivist countries, high MAS countries, low PDI countries, long term orientation countries, and high uncertainty avoidance countries, for applicable values. Orange represents individualisic countries, low MAS countries, high PDI countries, short term orientation countries, and low uncertainty avoidance countries, for applicable values (continuation).
Table 10: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs.
Table 11: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
Table 12: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
Table 13: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
Table 14: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs (continuation).
Table 15: Graphs showing correlations between percentage of responses indicating a value and the country’s value across all approaches, values, and LLMs.